# MimicKit: A Reinforcement Learning Framework for Motion Imitation and Control

### v2

@@ 3.4. Engine
While the Environment implements the high-level logic for simulating a particular task, the low-level simulation of the world is delegated to an Engine. The Engine class, implemented in mimickit/engines/engine.py, provides a unified API that abstracts away the low-level details of how an Environment is simulated. Different Engines can be constructed for different physics engines and real-world robotic systems. This enables a specific task and environment to be instantiated through different underlying simulators and physical robots. MimicKit currently only supports IsaacGym (Makoviychuk et al., 2021), but additional Engines will be introduced in the future to support other physics simulators, as well as deployment on real robots.
@@ 5.1. Installation
MimicKit can be installed by following the steps below:
@@
    MimicKit utilizes NVIDIA IsaacGym for high-performance physics simulation. IsaacGym installation instructions can be found at: https://developer.nvidia.com/isaac-gym.
  2. (2)
    Next install the dependencies from requirements.txt:
@@ 5.2. Training
--visualize true --log_file output/log.txt \
--out_model_file output/model.pt
@@
  • --log_file specifies the output log file, which will keep track of statistics during training.
  • --out_model_file specifies the output model file, which contains the model parameters.
@@ 5.3. Distributed Training
--num_workers 2 --device cuda:0
where --num_workers specifies the number of worker processes used to parallelize training. --device specifies the device used for training, which can be either cpu or cuda:0. When training with multiple GPUs, the number of worker processes used to parallelize training must be less than or equal to the number of GPUs available on the system.

### v3

@@ 3.4. Engine
While the Environment implements the high-level logic for simulating a particular task, the low-level simulation of the world is delegated to an Engine. The Engine class, implemented in mimickit/engines/engine.py, provides a unified API that abstracts away the low-level details of how an Environment is simulated. Different Engines can be constructed for different physics engines and real-world robotic systems. This enables a specific task and environment to be instantiated through different underlying simulators and physical robots. MimicKit currently supports Isaac Gym (Makoviychuk et al., 2021), Isaac Lab (NVIDIA et al., 2025), and Newton (Newton Contributors, 2025). Additional Engines will be introduced in the future to support other physics simulators, as well as deployment on real robots.
@@ 5.1. Installation
MimicKit supports different simulator backends. First, install the simulator of your choice from their respective source. The simulators that are currently supported include: Isaac Gym (Makoviychuk et al., 2021), Isaac Lab (NVIDIA et al., 2025), and Newton (Newton Contributors, 2025). MimicKit can then be installed by following the steps below:
@@
    Install the dependencies from requirements.txt:
@@ 5.2. Training
--engine_config data/engines/isaac_gym_engine.yaml \
@@
--visualize true \
--out_dir output/
@@
  • --engine_config specifies the configuration file for the Engine to select different simulator backends.
@@
  • --log_dir specifies the output directory where the models and logs will be saved.
@@ 5.3. Distributed Training
--devices cuda:0 cuda:1
where --devices specifies the devices used for training, which can be cpu or cuda:{i}. Multiple devices can be provided to parallelize training across multiple processes.
@@ 5.4. Testing
5.5. Visualizer User Interface
A simple interface is provided for controlling the visualizer, which is shared across different Engines.
  • Camera control: Hold Alt key and drag with the left mouse button to pan the camera. Scroll with the mouse wheel to zoom in/out.
  • Pause Simulation: Enter key can be used to pause/unpause the simulation
  • Step Simulation: Space key can be used to step the simulator one step at a time.
Each Engine may also implement different additional controls depending on the underlying simulator.